OpenStudy (goldrush18):

The roots of the equation x^2+4x+1=0 are alpha and beta. Without solving the equation

OpenStudy (goldrush18):

i. state the values of \[\alpha+ \beta and \alpha \beta\]

OpenStudy (goldrush18):

ii. find the value of\[\alpha^2+\beta^2\]

OpenStudy (goldrush18):

iii. find the equation whose roots are \[1 + 1 \div \alpha and 1+1 \div \beta\]

OpenStudy (maheshmeghwal9):

\[\alpha+\beta=-b/a=-4/1\]

OpenStudy (maheshmeghwal9):

\[\alpha \beta=c/a=1/1=1.\]

OpenStudy (maheshmeghwal9):

\[\alpha^2+\beta^2=(\alpha+\beta)^2-2\alpha \beta\] put the values

OpenStudy (goldrush18):

ok

OpenStudy (goldrush18):

how do i go about doing the third part

OpenStudy (maheshmeghwal9):

u put the roots as\[x^2-(\gamma+\delta)x+\gamma \delta=0\] where \[\delta and \gamma\] are the 3rd part's roots.

OpenStudy (goldrush18):

im a bit confused

OpenStudy (maheshmeghwal9):

what type of conf.

OpenStudy (goldrush18):

oh now i go about the third part that says find the equation because im trying to do it but not sure if its correct can i tell u the answer what i get when im finished